# A Survey of Controllable Learning: Methods and Applications in Information Retrieval

### v2

@@ 4.1.1 Rule-Based Techniques
Rule-based controllable techniques have proven to be indispensable in many scenarios such as recommender systems, information retrieval, and text generation due to its direct mode. Following Definition 2.1, the $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents the system performance we expect to achieve, such as security, fairness etc and the $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ is always implicit. Rule-based methods acting as a patchwork solution, enhance system performance from a product perspective. Specifically, rule-based methods can be categorized into two types: pre-processing and post-processing. Formally, the control function $h$ is defined as a mapping: $h:f\rightarrow f\circ g_{\text{rule}}$ in the context of pre-processing, whereas it is defined as: $h:f\rightarrow g_{\text{rule}}\circ f$ in the context of post-processing, where $f$ represents the base function and $g_{\text{rule}}$ denotes the rule-based control mechanism. For pre-processing algorithms, the strategy involves handling the context information including user profiles, interaction history, and others, denoted as $\bm{s}_{\mathrm{ctx}}$ in Definition 2.1, with the aim of altering the representations of users and items to meet specific requirements, such as privacy preservation or fairness. In contrast, post-processing algorithms, such as removing outdated items or promoting less popular items, can directly enhance diversity.
@@ 4.1.2 Pareto Optimization
In the domain of machine learning, the concept of controllability has garnered significant attention due to the necessity to balance multiple, often conflicting objectives inherent in lots kinds of tasks. The integration of Pareto optimality principles from multi-objective optimization (MOO) has proven particularly instrumental in crafting algorithms that can navigate these trade-offs effectively. Following Definition 2.1, the $\bm{s}_{\mathrm{tgt}}$ of the triplet typically represents a single Pareto optimal solution that simultaneously satisfies multiple target objectives. The $\bm{s}_{\mathrm{desc}}$ generally consists of a set of multi-objective weights or constraints that enable the system to achieve the target. The $\bm{s}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context. As illustrated in Figure 7, the key challenge of Pareto optimization in the domain of controllable learning is determining how to guide the learner $f$ to achieve Pareto optimality under given constraints.
@@ 4.1.3 Hypernetwork
Hypernetworks [83], which are neural networks designed to generate the parameters for another network, offer a flexible and efficient way to manage and adapt model parameters dynamically. Therefore, they have gradually emerged as a key technique for enhancing model controllability due to their explicit control capabilities. Following Definition 2.1, the $\bm{s}_{\mathrm{desc}}$ typically represents the description of the task or domain, usually serving as the input for the hypernetwork. The hypernetwork act as the control function $h$ taking $\bm{s}_{\mathrm{desc}}$ as input, mapping the learner through parameter generation. Thus the mapped learner could achieve the corresponding $\bm{s}_{\mathrm{tgt}}$.
@@
Galanti and Wolf [37] discuss the modularity of hypernetworks, emphasizing their ability to decompose a complex task into smaller, manageable sub-tasks. This modular approach allows for the generation of highly specialized parameters tailored to specific retrieval tasks, thereby improving accuracy and efficiency. Hyperprompt [36] introduces a prompt-based task-conditioning method for transformers that leverages hypernetworks. This approach dynamically generates prompts that condition the transformer model to adapt to different retrieval tasks, enabling more precise and context-aware information retrieval. Hamur [13] propose a hyper adapter designed for multi-domain recommendation. Hamur utilizes hypernetworks to adaptively generate domain-specific parameters, thereby enhancing the model’s ability to recommend items across different domains with high relevance and accuracy. Furthermore, CMR [3] explore the use of policy hypernetworks for controllable multi-objective re-ranking in recommender systems. By leveraging hypernetworks to generate parameters that balance multiple objectives, such as relevance and diversity, the system can dynamically adjust its recommendations based on user preferences and system goals. Similarly, HyperBandit  [12] maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. Mahabadi et al. [84] discuss parameter-efficient multi-task fine-tuning for transformers via shared hypernetworks, allowing for the efficient sharing of parameters across multiple tasks, reducing the overall computational load and improving the scalability of transformer models.
@@ 4.1.4 Other Methods
There are also some control methods in machine learning, which is not strictly belong to the above three classifications. Some studies [85, 16, 15, 86] propose the concept of Disentanglement , which decouples user interests into specific dimensions within the latent space to facilitate control. Specifically, Multi-VAE and Supervised $\beta$-VAE [85, 15] provide the ‘knobs’ for users, with each knob corresponding to an item aspect. The latent space where generative factors (here, a preference towards an item category like genre) are captured independently in their respective dimensions, thereby enabling predictable manipulations. LP [16] models user-item interactions and user requirements in a linear space, enabling controllable user updates by integrating user requirements into user features through a weighted summation approach. CGIR [86] propose a weakly-supervised method that can learn a disentangled item representation from user-item interaction data and ground the semantic meaning of attributes to dimensions of the item representation. During inference, CGIR start from the reference item and “walk” along the direction of the modification in the item representation space to retrieve a sequence of items in a gradient manner. Test-Time Adaption (TTA) means the algorithms do not change the model through retraining manner. especially in the scenario of online recommendation, where time consuming is quite strict, some works focus on how to change the model to adapt the dynamic environment directly in test stage. Reinforcement Learning (RL) is a kind of classic method to achieve controllability. Since these algorithms learn from the interaction with the environment (actually the reward function of the environment), it is possible for us to control the algorithms through dedicatedly designing the reward function. LangPTune [17] leverages reinforcement learning to enable the encoder to better encode implicit user interaction sequences into interpretable text. Similarly, Lu et al. [5] adopt a reinforcement learning approach to design fine-grained reward functions for control rules, allowing the model to accurately respond to specific control instructions. Additionally, inspired by the exceptional general intelligence of Large Language Models (LLMs) [63], researchers have begun to explore the LLMs application in recommender systems and other fields. RecLM-gen [5] constructs Supervised Fine-Tuning (SFT) tasks, augmented with labels derived from a conventional recommender model, explicitly improving LLMs’ proficiency in adhering to recommendation specific instruction which could be regarded as task description $\bm{s}_{\mathrm{desc}}$ as defined in Sec.2. For example, DLCRec [87] is a novel framework designed to enhance diversity control in LLM-based recommender systems. By decomposing the recommendation task into three supervised fine-tuning tasks-genre prediction, genre filling, and item prediction—it enables fine-grained diversity control guided by user-defined signals. Apart from the aforementioned technical approaches, recent works have leveraged large language models (LLMs) to enable interpretable and controllable Natural Language Interest Editing (NLIE). Notable examples include LACE [6], TEARS [9], and LangPTune [17], among others.
@@
HyperBandit [12] 2023 scenario adaptation control platform-mediated control hypernetwork in-processing
@@ 4.2.1 Pre-Processing Methods
Mysore et al. [6] introduce the LACE model, which crafts user profiles using intelligible concepts extracted from user-document interactions. This innovation permits users to modify their profiles, thereby directly influencing the recommendation outcomes. Through comprehensive offline assessments and user studies, LACE’s capacity to refine recommendation quality via user interaction has been convincingly validated. Lastly, Wang et al. [7] tackle the pivotal concerns of user privacy and the degree of control over data utilization in model training. This framework grants users the discretion to decide which of their interactions may be harnessed for training purposes, delicately balancing between optimizing recommendation effectiveness and honoring user preferences. The implementation of an influence function-based model, alongside an augmented model featuring multiple anchor actions, evidences the feasibility of reconciling high-quality recommendations with user consent. UCRS [2] imagines a counterfactual world where out-of-date user representations are discarded, and estimates their effects as the difference between factual and counterfactual worlds. After deducting such effects, incorporates the control command into recommender inference. As to user-feature controls, it revises the user feature specified by the control command (e.g., changing age from middle age to teenager) to conduct the final inference at the two levels. As to item-feature controls, UCRS adopts a user-controllable ranking policy to control the recommendations w.r.t. item category.

### v3

@@ 4.1.1 Rule-Based Techniques
Rule-based controllable techniques have proven to be indispensable in many scenarios such as recommender systems, information retrieval, and text generation due to its direct mode. Following Definition 2.1, the $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents the system performance we expect to achieve, such as security, fairness etc and the $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ is always implicit. Rule-based methods acting as a patchwork solution, enhance system performance from a product perspective. Specifically, rule-based methods can be categorized into two types: pre-processing and post-processing. Formally, the control function $h$ is defined as a mapping: $h:f\rightarrow f\circ g_{\text{rule}}$ in the context of pre-processing, whereas it is defined as: $h:f\rightarrow g_{\text{rule}}\circ f$ in the context of post-processing, where $f$ represents the base function and $g_{\text{rule}}$ denotes the rule-based control mechanism. For pre-processing algorithms, the strategy involves handling the context information including user profiles, interaction history, and others, denoted as $\bm{s}_{\mathrm{ctx}}$ in Definition 2.1, with the aim of altering the representations of users and items to meet specific requirements, such as privacy preservation or fairness. In contrast, post-processing algorithms, such as removing outdated items or promoting less popular items, can directly enhance diversity. For example, on an e-commerce platform, a user may not want the recommendation system to perceive their preference for a certain category of products (e.g., due to bias introduced by accidental clicks), which represent the task target $\bm{s}_{\mathrm{tgt}}$. A rule-based controllable recommendation system in the pre-processing stage should allow the user to specify a control mechanism $g\_{\text{rule}}$ to process the input to the recommendation system (i.e., the user’s interaction history), filtering out unintended user behaviors. Thus, it avoids modeling biased information and improves recommendation accuracy.
@@ 4.1.2 Pareto Optimization
In the domain of machine learning, the concept of controllability has garnered significant attention due to the necessity to balance multiple, often conflicting objectives inherent in lots kinds of tasks. The integration of Pareto optimality principles from multi-objective optimization (MOO) has proven particularly instrumental in crafting algorithms that can navigate these trade-offs effectively. Following Definition 2.1, the $\bm{s}_{\mathrm{tgt}}$ of the triplet typically represents a single Pareto optimal solution that simultaneously satisfies multiple target objectives. The $\bm{s}_{\mathrm{desc}}$ generally consists of a set of multi-objective weights or constraints that enable the system to achieve the target. The $\bm{s}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context. As illustrated in Figure 7, the key challenge of Pareto optimization in the domain of CL is determining how to guide the learner $f$ to achieve Pareto optimality under given constraints.
@@ 4.1.3 Hypernetwork
Hypernetworks [97], which are neural networks designed to generate the parameters for another network, offer a flexible and efficient way to manage and adapt model parameters dynamically. Therefore, they have gradually emerged as a key technique for enhancing model controllability due to their explicit control capabilities. Following Definition 2.1, the $\bm{s}_{\mathrm{desc}}$ typically represents the description of the task or domain, usually serving as the input for the hypernetwork. The hypernetwork act as the control function $h$ taking $\bm{s}_{\mathrm{desc}}$ as input, mapping the learner through parameter generation. Thus the mapped learner could achieve the corresponding $\bm{s}_{\mathrm{tgt}}$. Take the streaming recommendation as an example. It is known that data undergoes distributional shifts over time (e.g., users may prefer anime on weekends and educational videos during weekdays). Traditional train-test two-stage models fail to generalize to out-of-distribution data at test time. A feasible solution is to use a hypernetwork as the control function $h$, taking the time feature as the task description $\bm{s}_{\mathrm{desc}}$. By feeding $\bm{s}\_{\mathrm{desc}}$ into the hypernetwork, it outputs all or part of the weights for the learner $f$, thereby enabling the customization of a specific model for different time periods during the test stage and allowing the model to adapt to distributional shifts in the data.
@@
Galanti and Wolf [41] discuss the modularity of hypernetworks, emphasizing their ability to decompose a complex task into smaller, manageable sub-tasks. This modular approach allows for the generation of highly specialized parameters tailored to specific retrieval tasks, thereby improving accuracy and efficiency. Hyperprompt [39] introduces a prompt-based task-conditioning method for transformers that leverages hypernetworks. This approach dynamically generates prompts that condition the transformer model to adapt to different retrieval tasks, enabling more precise and context-aware information retrieval. Hamur [13] propose a hyper adapter designed for multi-domain recommendation. Hamur utilizes hypernetworks to adaptively generate domain-specific parameters, thereby enhancing the model’s ability to recommend items across different domains with high relevance and accuracy. Furthermore, CMR [3] explore the use of policy hypernetworks for controllable multi-objective re-ranking in recommender systems. By leveraging hypernetworks to generate parameters that balance multiple objectives, such as relevance and diversity, the system can dynamically adjust its recommendations based on user preferences and system goals. Similarly, HyperBandit  [12] maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. Mahabadi et al. [98] discuss parameter-efficient multi-task fine-tuning for transformers via shared hypernetworks, allowing for the efficient sharing of parameters across multiple tasks, reducing the overall computational load and improving the scalability of transformer models. Hypencoder [40] employs a hypernetwork to generate a dedicated lightweight neural network for each query, which is then used to score documents. This approach overcomes the representational limitations of traditional vector dot product methods and enables fine-grained control over the relevance evaluation function.
@@ 4.1.4 Other Methods
There are also some control methods in machine learning, which is not strictly belong to the above three classifications. Some studies [99, 16, 15, 100] propose the concept of Disentanglement , which decouples user interests into specific dimensions within the latent space to facilitate control. Specifically, Multi-VAE and Supervised $\beta$-VAE [99, 15] provide the ‘knobs’ for users, with each knob corresponding to an item aspect. The latent space where generative factors (here, a preference towards an item category like genre) are captured independently in their respective dimensions, thereby enabling predictable manipulations. LP [16] models user-item interactions and user requirements in a linear space, enabling controllable user updates by integrating user requirements into user features through a weighted summation approach. CGIR [100] propose a weakly-supervised method that can learn a disentangled item representation from user-item interaction data and ground the semantic meaning of attributes to dimensions of the item representation. During inference, CGIR start from the reference item and “walk” along the direction of the modification in the item representation space to retrieve a sequence of items in a gradient manner. Reinforcement Learning (RL) is a kind of classic method to achieve controllability. Since these algorithms learn from the interaction with the environment (actually the reward function of the environment), it is possible for us to control the algorithms through dedicatedly designing the reward function. LangPTune [17] leverages reinforcement learning to enable the encoder to better encode implicit user interaction sequences into interpretable text. Similarly, Lu et al. [5] adopt a reinforcement learning approach to design fine-grained reward functions for control rules, allowing the model to accurately respond to specific control instructions. Qin et al. [101] leveraged the reflective reasoning capabilities of LLMs to summarize user preferences from multiple perspectives and integrated reinforcement learning to dynamically adapt to changes in user preferences. Zhang et al. [102] propose the user-oriented exploration policy, which controls exploration patterns across user groups by customizing intensities via a distributional critic, quantile-targeting actors, and regularization for diversity and stability Additionally, inspired by the exceptional general intelligence of Large Language Models (LLMs) [73], researchers have begun to explore the LLMs application in recommender systems and other fields. GOMMIR [18] enhances user control over the recommendation process by integrating REINFORCE with both verbal and non-verbal user feedback. It further optimizes recommendation performance through goal-oriented rewards and hard negative sampling. Similarly, MocDT [103] leverages a Decision Transformer to represent and prioritize multiple objectives by introducing adjustable multi-objective control signals. This enables the generation of recommendation sequences tailored to different goals without the need for retraining, thereby enhancing the flexibility and controllability of recommendation systems in multi-objective scenarios. RecLM-gen [5] constructs Supervised Fine-Tuning (SFT) tasks, augmented with labels derived from a conventional recommender model, explicitly improving LLMs’ proficiency in adhering to recommendation specific instruction which could be regarded as task description $\bm{s}_{\mathrm{desc}}$ as defined in Sec.2. For example, DLCRec [104] is a novel framework designed to enhance diversity control in LLM-based recommender systems. By decomposing the recommendation task into three supervised fine-tuning tasks-genre prediction, genre filling, and item prediction—it enables fine-grained diversity control guided by user-defined signals. Apart from the aforementioned technical approaches, recent works have leveraged large language models (LLMs) to enable interpretable and controllable Natural Language Interest Editing (NLIE). Notable examples include LACE [6], TEARS [9], and LangPTune [17], among others. Test-Time Adaptation refers to algorithms that do not modify the model through retraining; instead, they allow a pre-trained model to leverage unlabeled data during the testing phase before making predictions. This type of algorithm is well-suited for online scenarios, where time constraints are stringent. Some studies focus on adapting the model directly during the test stage to cope with dynamic environments. $T^{2}ARec$ [105] innovatively integrates Test-Time Training into sequential recommendation by introducing two alignment-based self-supervised losses. HyperBandit [12] employs a hypernetwork to dynamically adjust the model based on temporal features of the test data before making predictions, without the need for retraining. Hamur [13] employs a domain-shared hypernetwork that implicitly captures shared information across domains and dynamically generates adapter parameters to adapt to domain shifts during the test stage. Notably, not all hypernetwork-based algorithms fall under this category. This is because hypernetworks are typically designed to support proactive control by users or platforms, whereas test-time adaptation generally requires reactive adjustments based on test data.
@@
MocDT [103] 2025 multi-objective control user-centric control RL in-processing
@@
FollowIR [45] 2024 user portrait control user-centric control SFT pre-processing
InstructIR [46] 2024 user portrait control user-centric control SFT pre-processing
@@
HyperBandit [12] 2023 scenario adaptation control platform-mediated control hypernetwork, TTA in-processing
@@ 4.2.1 Pre-Processing Methods
Mysore et al. [6] introduce the LACE model, which crafts user profiles using intelligible concepts extracted from user-document interactions. This innovation permits users to modify their profiles, thereby directly influencing the recommendation outcomes. Through comprehensive offline assessments and user studies, LACE’s capacity to refine recommendation quality via user interaction has been convincingly validated. Lastly, Wang et al. [7] tackle the pivotal concerns of user privacy and the degree of control over data utilization in model training. This framework grants users the discretion to decide which of their interactions may be harnessed for training purposes, delicately balancing between optimizing recommendation effectiveness and honoring user preferences. The implementation of an influence function-based model, alongside an augmented model featuring multiple anchor actions, evidences the feasibility of reconciling high-quality recommendations with user consent. UCRS [2] imagines a counterfactual world where out-of-date user representations are discarded, and estimates their effects as the difference between factual and counterfactual worlds. After deducting such effects, incorporates the control command into recommender inference. As to user-feature controls, it revises the user feature specified by the control command (e.g., changing age from middle age to teenager) to conduct the final inference at the two levels. As to item-feature controls, UCRS adopts a user-controllable ranking policy to control the recommendations w.r.t. item category. It is worth noting that, besides those works on recommender systems, the use of instructions in information retrieval is a relatively new development. TART [43] and Instructor [44] adopt simple task prefixes during training. Meanwhile, several benchmark efforts explicitly focus on evaluating the instruction-following ability of retrievers, such as FollowIR [45] and InstructIR [46].
@@ 8 Conclusion
Acknowledgement
This work was partially supported by the National Natural Science Foundation of China (No. 62376275, 62472426). Work partially done at Beijing Key Laboratory of Research on Large Models and Intelligent Governance, and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE. Supported by fund for building world-class universities (disciplines) of Renmin University of China.